OpenStudy (mikezack123):

product of (4x)(-3y) :) medal will be given. thanks

OpenStudy (anonymous):

do you know how to multiply?? :)

OpenStudy (anonymous):

\[ (4x)(-3y) =4 \times (-3) \times x \times y =-12xy\]

OpenStudy (mathstudent55):

Multiply the coefficients (the numbers) and multiply the variables (the letters). For example: (2a)(-5b) = -10ab